### v1

# Femtoscopy of $DN$ and $\bar{D}N$ systems

Abstract
The capability of the ALICE@LHC and STAR@RHIC experiments to reconstruct $D$ mesons has enabled femtoscopic correlation measurements of open-charm mesons in both small and large systems. In this work, we present a theoretical calculation of the correlation functions of $D$ and $\bar{D}$ mesons with nucleons, based on the Koonin-Pratt formalism applied to the systems produced in these high-energy collisions. We use an effective Lagrangian that models the interaction between charmed mesons and baryons. From the unitarized on-shell amplitudes, we extract the low-energy scattering parameters and compute the correlation functions using the Lednický-Lyuboshitz approximation. Additionally, we apply the TROY formalism to obtain the off-shell $T$-matrix in coupled channels and reconstruct the full wave function, which we incorporate into the Koonin-Pratt formula. We compare these two approaches and provide predictions for different correlated pairs, including the effects of the Coulomb interaction and thermal weights for the coupled channels. These results can be tested against current and future experimental data from ALICE and STAR collaborations in both proton-proton and heavy-ion collisions.
Femtoscopy, heavy mesons, effective field theories, $T$-matrix equation, relativistic scattering theory
I Introduction
Quantum Chromodynamics (QCD) is known to be the fundamental theory of the strong interaction. Many processes at high energies (or shorter distances than the size of the nucleon) can be determined using a perturbative approach involving quark and gluon interactions. However, the low-energy QCD regime is strongly coupled and the interactions among hadrons, the relevant degrees of freedom at this energy scale, are not well determined and usually difficult to analyze experimentally.
Over the past years, femtoscopy has turned out to be a promising tool to study the interactions among hadrons (see the review of Ref. [1]). Femtoscopy is based on determining hadron-hadron correlations in momentum space as the ratio between the distribution of relative momenta for pairs produced in the same collision and in different collisions. Since the pioneer works in [2, 3] and further developments of the STAR Collaboration [4, 5], the ALICE Collaboration has taken the leading role and determined the correlations for different hadron-hadron systems (see, for example, the works in [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]).
In the past years the ALICE Collaboration has started investigating hadron-hadron interactions in the charm sector using femtoscopic techniques. More precisely, the correlation functions of $D^{(\pm)}\pi^{(\pm)}$ and $D^{(\pm)}K^{(\pm)}$ [26] as well as $D^{-}p$ and $D^{+}\bar{p}$ [27] in high-multiplicity pp collisions at 13 TeV have been recently obtained, thus giving new insights into interactions involving with $D$ mesons.
Motivated by these recent experimental developments on femtoscopy, several theoretical analyses have been carried out in the charm sector. Some of them have focused on the correlation functions of exotic hidden charmed meson states generated dynamically via the scattering of open-charm mesons, such as $X(3872)$ [28, 29], $T_{cc}$ [28, 30, 31], $Z_{c}$(3900) and $Z_{cs}(3985)$ [32], as well as the correlation functions of $P_{c}(4440)$ and $P_{c}(4457)$ that emerge from the interaction of $D$ mesons with baryons [33]. Also, some recent theoretical studies have aimed at determining the correlation functions of open-charm mesons with light mesons [34, 35, 36, 37, 38], and the charmonium-nucleon [39] or even the $\alpha$-charmonium [40] correlation functions.
In view of the recent experimental results on femtoscopy for $D$ mesons with protons [27] and following our previous analysis of the $D\pi$ correlations functions [35], in the present work we focus on the determination of the correlation functions for $D^{+}p$, $D^{-}p$, $D^{0}p$ and $\bar{D}^{0}p$ systems. For that purpose we start from the interaction of $D$ and $\bar{D}$ mesons with nucleons obtained from the vector meson-exchange model of Ref. [41] in the zero-range approximation [42]. The correlation functions are then calculated using the off-shell $T$-matrix TROY approach recently developed in [35]. We also compare our results with those obtained within the Lednický-Lyuboshitz (LL) approximation in order to assess the differences and caveats of this approximation. Within the simple LL approximation, we employ different $D-$ and $\bar{D}-$nucleon interactions [41, 43, 44, 45, 46, 47, 48, 49], so as to determine the dependence of the $D^{+}p$, $D^{-}p$, $D^{0}p$ and $\bar{D}^{0}p$ correlation functions on the underlying interaction model.
In Sec. II we provide theoretical details of the interaction model between charmed mesons and nucleons, the unitarization techniques to obtain the $T$-matrices and the wave functions, and the femtoscopy correlation functions. In Sec. III we present the results of the correlation functions using the LL approximation as well as the off-shell approach using the TROY formalism. In Sec. IV we give a summary of this work.
II Formalism
In this section, we provide theoretical details to compute the femtoscopy correlation functions for $DN$ and $\bar{D}N$ pairs. In Sec. II.1 we present the model that describes the strong interaction between the hadrons. In Sec. II.2 we provide details on how the $T$-matrix equation is solved both in the off-shell case and in the on-shell factorization approach. In Sec. II.3, we introduce the Koonin-Pratt (KP) equation as well as the two approaches we consider: the LL approximation based on the effective-range expansion of the on-shell amplitudes and the off-shell calculation within TROY.
II.1 Meson-baryon strong interaction
An effective description of meson-baryon scattering involving charm and anticharm degrees of freedom is developed in Ref. [41]. Using that approach, we focus on the $C=1$ and $C=-1$ sectors, to accommodate the $DN$ and $\bar{D}N$ channels, respectively. We follow the notation and conventions given in Ref. [42].
The lowest-order tree-level diagrams of the meson–baryon interaction are composed by vector meson exchanges [42]. The $s$-wave amplitude $V_{ij}$ (where $i$ and $j$ refer to the incoming and outgoing meson-baryon channels) is dominated by the $l=0$ projection of the $t$-channel meson exchange, while the $u$- and $s$-channel contributions only become more relevant in the $p$-wave amplitude, potentially playing a significant role at higher energies. In fact, in Ref. [50] these contributions were evaluated for the $\bar{K}N$ system and were found to amount to approximately 20% of the $t$-channel contribution at energies around 200 MeV above threshold. In the charm sector, such contributions are expected to be even smaller, as discussed in Ref. [42].
This $t$-channel vector meson exchange contribution can be computed using the Lagrangian defined in Ref. [43, 41], including all allowed vector-meson exchanges in a given transition $i\to j$. From Ref. [42], the $t$-channel exchanged potential reads,
$V^{ISC}_{ij}(q_{i},p_{i},q_{j},p_{j})=\frac{g^{2}}{4}\bar{u}(p_{j})\sum_{v}\frac{C^{ISC}_{ij,v}}{t-m_{v}^{2}}\left[\not{q}_{i}+\not{q}_{j}-\frac{q_{i}^{2}-q_{j}^{2}}{m_{v}^{2}}(\not{q}_{i}-\not{q}_{j})\right]u(p_{i})\ ,$ (1)
where $g$ is the universal vector meson coupling constant, $C^{ISC}_{ij,v}$ represents the isospin coefficients depending on the ($ISC$) quantum numbers as well as the $i$ and $j$ channels and the vector meson exchanged $v$ (see [41] for explicit values). The 4-momenta $q_{i}$ and $q_{j}$ correspond to the incident and outgoing mesons. Finally, $u(p_{i}),\bar{u}(p_{j}),p_{i},p_{j}$ are the incident and outgoing baryon wave functions, as well as their respective momenta.
Then, we assume the $|t|\ll m_{v}^{2}$ limit to simplify the $t$-dependence in the denominator as well as the third term in the brackets. This is equivalent to making a contact interaction at the vertex. As explained in Ref. [42], the error introduced in this approximation is small. Finally, we take the $s$-wave projection of this so-called Zero-Range (ZR) approximation for the interaction kernel,
$V_{ij,\ \textrm{ZR}}^{ISC}(|\bm{q}_{i}|=|\bm{p}_{i}|,|\bm{q}_{j}|=|\bm{p}_{j}|;\sqrt{s})=-\dfrac{N}{4}g^{2}\sum_{v}\dfrac{C^{ISC}_{ij,v}}{m_{v}^{2}}\left(2\sqrt{s}-E_{i}-E_{j}+\frac{p_{i}^{2}}{E_{i}+M_{i}}+\frac{p_{j}^{2}}{E_{j}+M_{j}}\right)\ ,$ (2)
where, in the center of mass (CM) frame,
$N=\sqrt{\dfrac{M_{i}+E_{i}(\bm{p}_{i})}{2M_{i}}}\ \sqrt{\dfrac{M_{j}+E_{j}(\bm{p}_{j})}{2M_{j}}}\ ,$ (3)
is a normalization factor arising from the Dirac spinors; and $M_{i}$, $M_{j}$ and $E_{i}$, $E_{j}$ are the masses and energies of the baryons in channels $i$ and $j$, respectively. Equation (2) serves as an off-shell kernel for the meson-baryon scattering in the ZR approximation. The on-shell version of [42] is simply achieved by setting $p_{i/j}^{2}=E_{i/j}^{2}-M_{i/j}^{2}$. In this work, both off-shell and on-shell versions of this potential in the ZR approximation will be used.
Finally, by assuming a common vector-meson mass, one can apply the Kawarabayashi-Suzuki-Fayyazuddin-Riazuddin rule [51, 52], $g=m_{v}/(\sqrt{2}f)$, where $f$ is the meson weak decay constant. In this case, the potential takes the form of the Weinberg–Tomozawa (WT) interaction,
$V_{ij,\ \textrm{WT}}^{ISC}(\sqrt{s})=-\dfrac{N}{8f^{2}}C_{ij}^{\textrm{WT}}\ (2\sqrt{s}-M_{i}-M_{j})\ ,$ (4)
where the new isospin coefficients
$C_{ij}^{\textrm{WT}}=\kappa\sum_{v}C^{ISC}_{ij,v}\ ,$ (5)
with a correction factor $\kappa=1/4$ only when the exchanged vector meson contains a charm quark, to balance from the universal vector mass assumption [42].
In this work, we focus on four isospin/strangess/charm sectors. To account for $D^{0}p$ and $D^{+}p$ cases we will consider the sector $\{ISC\}=\{0,0,1\}$ with 7 coupled channels,
$\pi\Sigma_{c},DN,\eta\Lambda_{c},K\Xi_{c},K\Xi^{\prime}_{c},D_{s}\Lambda,\eta^{\prime}\Lambda_{c}\ ,$ (6)
leaving aside heavier channels with additional $c\bar{c}$ content, and also the sector $\{ISC\}=\{1,0,1\}$ with 8 coupled channels,
$\pi\Lambda_{c},\pi\Sigma_{c},DN,K\Xi_{c},\eta\Sigma_{c},K\Xi^{\prime}_{c},D_{s}\Sigma,\eta^{\prime}\Sigma_{c}\ .$ (7)
For the $D^{-}p$ and $\bar{D}^{0}p$ we work with the $\{ISC\}=\{0,0,-1\}$ and $\{ISC\}=\{1,0,-1\}$ sectors, each of them with a single channel, $\bar{D}N$.
We anticipate that to compute the correlation functions with TROY, it is more convenient to work in the physical basis (with the physical mass of each state). In this case, the number of coupled channels is 16, 9, 2 and 1, for the cases $D^{0}p,D^{+}p,D^{-}p,\bar{D}^{0}p$, respectively.
II.2 Off- and on-shell $T$-matrices
The perturbative potential $V_{ij}(q^{\prime},q;\sqrt{s})$ is used as the kernel of the Bethe-Salpeter (or $T$-matrix) equation,
$T_{if}(q^{\prime},q;\sqrt{s})=V_{if}(q^{\prime},q;\sqrt{s})+\sum_{l}\int\frac{d^{3}k}{(2\pi)^{3}}\frac{2M_{l}(E_{l}+\omega_{l})}{2E_{l}\omega_{l}}\frac{V_{il}(q^{\prime},k;\sqrt{s})T_{lf}(k,q;\sqrt{s})}{s-(E_{l}+\omega_{l})^{2}+\mathrm{i}\eta}\ ,$ (8)
where we denote the initial and final CM momenta as $q^{\prime}=|\bm{q}_{i}|$ and $q=|\bm{q}_{f}|$, respectively. The quantities $E_{l}=\displaystyle\sqrt{M_{l}^{2}+k^{2}}$ and $\omega_{l}=\displaystyle\sqrt{m_{l}^{2}+k^{2}}$ are the intermediate baryon and meson energies, respectively. The on-shell CM momentum of the final ($f$) baryon-meson pair is $q=\lambda^{1/2}(s,M_{f}^{2},m_{f}^{2})/(2\sqrt{s})$, where $\lambda$ is the Källén function. The initial CM momentum of the pair $q^{\prime}$ is not fixed by $\sqrt{s}$ since in Eq. (8) we consider the half off-shell $T$-matrix elements.
To regularize the ultraviolet (UV) limit of the perturbative potential in Eq. (8) we apply, as in Ref. [35], an exponential form factor,
$F(q^{\prime},q)=\exp\left(-\frac{q^{\prime 2}+q^{2}}{\Lambda^{2}}\right),$ (9)
to the potential, with the regulator $\Lambda$, universal for all scattering channels, to be fixed as we will explain later.
In the on-shell approximation, $q^{\prime}$ is taken to be $q^{\prime}=\lambda^{1/2}(s,M_{i}^{2},m_{i}^{2})/(2\sqrt{s})$ and both $V_{il},T_{lf}$ can be pulled out of the integral. One gets
$T_{if}(\sqrt{s})=V_{if}(\sqrt{s})+\sum_{l}V_{il}(\sqrt{s})G_{l}(\sqrt{s})T_{lf}(\sqrt{s})\ ,$ (10)
with
$G_{l}(\sqrt{s})=2M_{l}\int\frac{d^{3}k}{(2\pi)^{3}}\frac{E_{l}+\omega_{l}}{2E_{l}\omega_{l}}\frac{1}{s-(E_{l}+\omega_{l})^{2}+\mathrm{i}\eta}\ .$ (11)
In the on-shell version, the loop function $G_{l}(\sqrt{s})$ also needs to be regularized, which we will do by setting a UV 3-momentum cut-off, $k_{\textrm{max}}$ and using the analytical expression given in Ref. [53]. We will also use a universal value of $k_{\textrm{max}}$ for all $DN$ and $\bar{D}N$ channels.
II.3 Femtoscopy correlation function
The primary quantity observed in femtoscopy is the correlation function of a given channel $f$. It is expressed as a function of the reduced relative momentum of two particles in their rest frame, determined using the KP formula [54, 55],
$C(k)=\int d^{3}r\ S(r)|\Psi_{f}(\bm{k};\bm{r})|^{2}\ ,$ (12)
where $S(r)$ is the source function, describing the spatial emission of particles, for which we assume a Gaussian profile,
$S(r)=(4\pi r_{0}^{2})^{-3/2}\ \exp\left(-\dfrac{r^{2}}{4r_{0}^{2}}\right)\ ,$ (13)
where $r_{0}$ defines the size of the source. In Eq. (12), $\Psi_{f}(\bm{k};\bm{r})$ is the wave function of the two-particle system, which encodes information about their interaction. It can be addressed by solving the appropriate Schrödinger equation or, as we do in this work, using the Lippmann-Schwinger equation after solving the $T$-matrix equation (8).
II.3.1 LednickĂ˝-Lyuboshitz approximation
A simple approximation for computing the correlation function is given in Ref. [56], where the wave function is taken to be the asymptotic one, even for $r\rightarrow 0$, with a combination of the incident plane wave plus the spherical outgoing wave modulated by the scattering amplitude $f(k)$. The latter is usually taken in the effective-range expansion, whose inverse follows an expansion in powers of the relative momentum $k$. When the Coulomb effects are included, the correlation function in the LL model reads [57, 29],
$C^{\textrm{C}}_{\textrm{LL}}(k)=A_{\textrm{C}}(k)\left\{1+\dfrac{|f_{\textrm{C}}(q)|^{2}}{4r_{0}^{2}}\left[1+e^{-(2kr_{0})^{2}}+A_{\textrm{C}}^{2}(1-e^{-(2kr_{0})^{2}})-\dfrac{d_{0}}{\sqrt{\pi}r_{0}}\right]\dfrac{2\ {\rm Re}f_{\textrm{C}}(k)}{\sqrt{\pi}r_{0}}F_{1}(2kr_{0})-\dfrac{A_{\textrm{C}}\ {\rm Im}f_{\textrm{C}}(k)}{r_{0}}F_{2}(2kr_{0})\right\}\ ,$ (14)
where $F_{1}(x)$ and $F_{2}(x)$ are functions defined as,
[…]
The functions $A_{\textrm{C}}(q)$ and $f_{\textrm{C}}(q)$ are the Gamow factor and the scattering amplitude with Coulomb effects, respectively,
$A_{\textrm{C}}(k)$ $=\dfrac{2\pi\gamma}{{\rm{exp}}\left(2\pi\gamma\right)-1}\ ,$ (16)
$f_{\textrm{C}}(k)$ $=\left(\dfrac{1}{a_{0}}+\dfrac{1}{2}d_{0}k^{2}-2k\gamma h(\gamma^{-1})-ikA_{\textrm{C}}(k)\right)^{-1}\ ,$ (17)
where $\gamma=Z_{1}Z_{2}\mu\alpha/k$ is the Sommerfeld factor, and $a_{0}$ and $d_{0}$, are the scattering length and effective range, respectively. The scattering length and effective range will be obtained from the on-shell $T$ matrix. In our convention, $a_{0}>0$ indicates attraction, while $a_{0}<0$ indicates repulsion or the presence of a bound state. Finally, the function $h(x)$ is defined as,
$h(\gamma^{-1})=-\log(|\gamma|)+\frac{1}{2}\psi(1-i\gamma)+\frac{1}{2}\psi(1+i\gamma)\ ,$ (18)
where $\psi(z)$ is the digamma function. In the absence of Coulomb interaction, $\gamma\rightarrow 0$, we recover the standard LL correlation function [58],
$C_{\textrm{LL}}(k)$ $=1+\dfrac{1}{2}\left|\dfrac{f(k)}{r_{0}}\right|^{2}\left(1-\dfrac{d_{0}}{2\sqrt{\pi}r_{0}}\right)$
$+\dfrac{2{\rm Re}f(k)}{\sqrt{\pi}r_{0}}F_{1}(2kr_{0})-\dfrac{{\rm Im}f(k)}{r_{0}}F_{2}(2kr_{0})\ ,$ (19)
where $f(k)$ is the $\gamma=0$ limit of $f_{\textrm{C}}(k)$.
It is important to note that the LL approximation does not account for the effects of coupled channels. Although the computation of the $T$ matrix does include them, the scattering parameters used in the formulas are obtained only for the direct (elastic) channel.
II.3.2 Off-shell $T$-matrix (TROY) approach
The TROY (T-matrix-based Routine for hadrOn femtoscopY) formalism uses the solution of the half off-shell $T$-matrix equation (8) to construct the scattering pair wave function [35].
Since the $T$-matrix equation contains all channels coupled to the one that is observed, one needs a more general KP formula,
$C(k)=\int d^{3}r\sum_{i}w_{i}\ S(r)|\Psi_{if}(\bm{k};\bm{r})|^{2}\ ,$ (20)
where the weights $w_{i}$ measure the production probability of the different pair combinations at freeze-out.
The calculation of $w_{i}$ requires a model for thermal production of hadronic states, plus feed-down corrections for those states that populate the final hadrons beyond the freeze-out. Therefore, it can be based on a statistical thermal model, such as Thermal-FIST [59], and needs information of the type of collision at hand (centrality, collision energy, kinematic cuts of the detected particles, efficiency of measured yields, etc.). It also needs to know the kinematic distribution of the particles that can lead to a pair correlation with a given relative momentum, see e.g. [60].
Leaving for a future study a more complete model for $w_{i}$, we adopt here a static fireball model and compute the thermal multiplicity of each species as [61]
$\frac{dn}{m_{\perp}dm_{\perp}dyd\varphi}\sim m_{\perp}\cosh{y}\exp\left(-m_{\perp}\cosh y/T\right)\ ,$ (21)
where $m_{\perp}=\sqrt{m^{2}+p_{\perp}^{2}}$, $m$ is the mass of the particle, $p_{\perp}$ its transverse momentum, $y$ is the rapidity, $\varphi$ is the azimuthal angle of the particle’s momentum, and $T$ is the temperature. We integrate this multiplicity without kinematic restrictions and use $T=171$ MeV for high-multiplicity pp collisions at $\sqrt{s}=13$ TeV at the LHC [22]. Then, the thermal weight $w_{i}$ for a given pair—using $w_{f}=1$ as the reference value—is given by
$w_{i}=\frac{n_{1,i}\times n_{2,i}}{n_{1,f}\times n_{2,f}}\ ,$ (22)
where $1,2$ denote the first and second particles of the pair. This prescription, while simple, is able to provide the correct order of magnitude of the weights. In particular, the pairs affected by feed down can deviate more from the actual weight; however, this is a better approach than setting to unity all weights. We have checked that the weights $w_{i}$ obtained by our prescription do not deviate much from those of the recent work [60] for the particular systems explored there.
Turning back to Eq. (20), the total wave function is splitted as,
$\Psi_{if}(\bm{k};\bm{r})=\delta_{if}[\Psi^{\textrm{C}}_{if}(\bm{k};\bm{r})-\Psi^{\textrm{C}}_{0,if}(kr)]+\varphi_{if}(k;r)\ ,$ (23)
where $\Psi^{\textrm{C}}_{if}(\bm{k};\bm{r})$ is the total Coulomb wave function (or plane wave in the absence of electromagnetic forces), $\Psi^{\textrm{C}}_{0,if}(kr)$ is its $s$-wave projection and $\varphi_{if}(k;r)$ is the $s$-wave projection of the total wave function (including strong and Coulomb interactions). In writing Eq. (23) we have assumed that at low energies only the $l=0$ wave is modified by the strong interaction. Specific details on the Coulomb wave functions are given in Ref. [35].
After solving Eq. (8), TROY computes the interacting wave function out of the off-shell $T$-matrix elements as,
$\varphi_{if}(k;r)$ $=j_{0}(kr)\delta_{if}$ (24)
+∫$+\int\frac{d^{3}q^{\prime}}{(2\pi)^{3}}\frac{2M_{i}(E_{i}+\omega_{i})}{2E_{i}\omega_{i}}\frac{T_{if}(q^{\prime},k;\sqrt{s})j_{0}(q^{\prime}r)}{s-(E_{i}+\omega_{i})^{2}+i\eta}\ ,$
where $j_{0}(x)$ is the spherical Bessel function of zeroth order, and $\sqrt{s}=\sqrt{M_{f}^{2}+k^{2}}+\sqrt{m_{f}^{2}+k^{2}}$.
III Results
In this section, we present the correlation function of $DN$ and $\bar{D}N$ pairs, obtained using different approximations. In Sec. III.1 we use the LL model with Coulomb interactions for small systems. In Sec. III.2 we apply the off-shell TROY formalism for the same small systems. In Sec. III.4 we use TROY to compare to ALICE results in pp collisions and predict the result for larger systems, as those studied by the STAR collaboration.
III.1 LL approximation
For the LL case, we apply the on-shell results by solving (10) using first the ZR approximation with $g=6.0$ and $k_{\textrm{max}}=674$ MeV. Then, for the WT interaction we use $f=1.15f_{\pi}$ with $f_{\pi}=92$ MeV and $k_{\textrm{max}}=752$ MeV.
First, we start by presenting the scattering lengths in the isospin basis for our model, while comparing our results with the outcome from other works. The scattering lengths are displayed in Tables 1 and 2 for $DN$ and $\bar{D}N$, respectively.
With regard to the $DN$ system, the $s$-wave $I=0$ and $I=1$ scattering lengths predicted by the WT model and those in the ZR approximation, both in the on-shell case, are very similar. This is an expected result, as the difference between the two approaches arises because of the slightly different energy-dependent $DN$ bare interaction. In both WT and ZR cases, the $I=0$ $\Lambda_{c}(2595)$ is generated dynamically at its mass by adjusting the cutoff, whereas in the $I=1$ sector a $\Sigma_{c}$ with a mass close to the $DN$ threshold of 2806 MeV is obtained, thus leading to a large scattering length. Note that the scattering lengths reported in Ref. [46] for the WT model of Ref. [44] slightly differ from those of the WT model shown here due to the slightly different cutoff used to regularize the loop function. Moreover, we show the scattering lengths of the $DN$ interaction of Ref. [41], reported in Ref. [43].
In that work, there is no $I=1$ quasi-bound state close to the $DN$ threshold but lies 180 MeV below, so the size of the corresponding scattering length is smaller than those from the WT and ZR approaches. As for the scattering lengths coming from the meson-exchange model of Ref. [46], they are qualitatively similar in both isospin sectors to those shown for WT and ZR models. This is due to the fact that the meson-exchange model also generates dynamically the $I=0$ $\Lambda_{c}(2595)$ and the $I=1$ $\Sigma_{c}(2800)$ for masses close to those obtained in our present study. And finally, for completeness, we show the scattering lengths for the SU(8) model of Ref. [45], reported in Ref. [46]. In that approach, both $I=0$ and $I=1$ scattering lengths are positive, being the $I=0$ one compatible with zero. The discrepancy with previous works has its origin in the different and richer resonant-hole composition close to $DN$ threshold, as described in Ref. [45].
In the case of $\bar{D}N$, the $I=0$ scattering length in the WT approach is zero, as the corresponding interaction is strictly vanishing in this channel. A result compatible with zero is consistently obtained in the ZR approach, as seen in Table 2. For $I=1$, both the WT and ZR prescriptions produce a similarly sizable negative scattering length, thus indicating a repulsive $\bar{D}N$ interaction in that sector. Upon comparing these results with those from Ref. [43], we observe that, while the $I=1$ value is very close to the ones for WT and ZR, there is a disagreement in the $I=0$ scattering length, that we are not able to trace back to the regularization method. In Ref. [47] a value close to zero was obtained for the $I=0$ scattering length, whereas a more repulsive scattering length than our results for WT and ZR was determined for $I=1$. This model is based on the exchange of mesons between $\bar{D}$ and $N$ supplemented by a short-range one-gluon exchange and, in fact, about half of the repulsive scattering length in $I=1$ comes from the hadronic meson-exchange contributions, which can be mapped to a certain extent to the WT interaction used in the present work. This result was revisited in Ref. [27], by modifying the coupling to the scalar meson, leading to a much more attractive interaction in $I=0$ and a value close to zero for $I=1$. Furthermore, the analysis of Ref. [49], with a model that considers a long range one-pion exchange interaction, obtains a bound $\bar{D}N$ state in $I=0$, leading to a very large and negative $I=0$ scattering length, whereas the $I=1$ one is much smaller. Notice that the value of Ref. [49] for the $I=0$ scattering length is quoted in [27] with an opposite sign, because the convention in the former paper is contrary to the ALICE one (and ours). The work of Ref. [48] uses a quark model that confines color and realizes dynamical chiral symmetry breaking, giving rise to values for $I=0$ and $I=1$ closer to the WT and ZR results, as reported in Ref. [27].
$DN$ $a_{0}~{}({I=0})[{\rm fm}]$ $a_{0}~{}(I=1)[{\rm fm}]$
[…]
Hofmann-Lutz [41, 43] −$-0.43$$-0.41$
[…]
Table 1: Scattering lengths for $DN$ with the Weinberg-Tomozawa (WT) and Zero-Range (ZR) approximations compared with other theoretical models in the literature.
$\bar{D}N$ $a_{0}~{}({I=0})[{\rm fm}]$ $a_{0}~{}(I=1)[{\rm fm}]$
WT 00 −$-0.29$
ZR $0.02$$-0.34$
Hofmann-Lutz [43] −$-0.16$$-0.26$
Haidenbauer [47] $0.07$$-0.45$
Haidenbauer,
$g_{\sigma}^{2}/4\pi=2.25$ [27] $0.67$ $0.04$
Yamaguchi [49, 27] −$-4.38$$-0.07$
Fontoura [48, 27] $0.16$$-0.25$
Table 2: Scattering length for $\bar{D}N$ with the Weinberg-Tomozawa and Zero-Range approximation compared with other theoretical models.
We now present the correlation functions for charged particles, obtained using the scattering lengths from this work and compare them with those obtained using the scattering lengths from previous studies, as discussed earlier. In all cases, we have used the LL approximation with Coulomb interaction of Eq. (14), employing a source radius of $r_{0}=1~{}\rm{fm}$, and a vanishing scattering range. We have checked in the $\bar{D}N$ case that the effect of the scattering range in the LL formula is negligible.
The correlation functions for the purely $I=1$ $D^{+}p$ system are shown in Fig. 1. The repulsive effect caused by the Coulomb interaction, leading to a correlation function below one, dominates the correlation function, although some differences can be appreciated between the models. The negative sign of the scattering length simulates a repulsive interaction in the majority of models, reducing further the correlation function with respect the the Coulomb only contribution. As mentioned before, the model of Ref. [45] presents a positive scattering length, hence an attractive interaction, leading to a higher correlation function compared to the other models and reaching values larger than one for a certain range of momenta.
The $D^{-}p$ correlation functions for different models are shown in Fig. 2. In this case, the attractive Coulomb effect dominates, leading in general to values of the correlation function larger than one. Since this system consists of a combination of isospins, the scattering length used is taken as $a=(a_{I=0}+a_{I=1})/2$ for each model. The variation in the scattering lengths presented earlier results in significantly different correlation functions. The small and negative averaged scattering lengths obtained in this work with the WT and ZR models lead to correlation functions that are very similar to those from other studies. The repulsive interaction associated with these values causes only a small deviation compared to the dominant attractive Coulomb effect. In contrast, the large negative scattering length reported in Ref. [49] for $I=0$ induces a substantial apparently repulsive effect, as seen in the dotted blue line in Fig. 2, where the correlation function is significantly lower than in the other works. This behavior is in fact signalling the presence of a bound state generated by the model of Ref. [49]. The opposite effect is observed in the results revisited in Ref. [27] using the model from Ref. [47]. In this case, the positive and large scattering length indicates an attractive interaction, leading to larger correlation function that agrees with experimental data [27].
Figure 1: $D^{+}p$ correlation function as function of the relative momentum. Our results with the ZR approximation are shown in black, in comparison with other works. All of them were calculated using the LL model, Eq. (14), with the low-energy scattering parameters from each paper.
Figure 2: $D^{-}p$ correlation function as function of the relative momentum. Our results with the ZR approximation are shown in black, in comparison with other works. All of them were calculated using the LL model, Eq. (14), with the low-energy scattering parameters from each paper.
III.2 TROY framework
In this subsection, we present the results obtained using the TROY formalism including all coupled channels in the off-shell $T$-matrix approach. We also use the ZR approximation with $g=6.0$ but adjust the value of the regulator to $\Lambda=925$ MeV so that the $I=0$ amplitude of the $DN$ channel has a maximum at the position of the $\Lambda_{c}(2595)$ state. Notice that the regulator $\Lambda$ plays an equivalent role to $k_{\textrm{max}}$ in the on-shell calculation, but they do not need to be equal since the off-shell calculation applies it via the form factor of Eq. (9), while the on-shell calculation uses a hard cutoff regularization.
Then, we apply the same regulator for the remaining ${\bar{D}}N$ and $DN$ channels. In particular, in the $I=1$ sector we find a maximum of the $DN$ scattering amplitude at $\sqrt{s}=2798$ MeV, which we identify with the $\Sigma_{c}(2800)$ [62], consistent with the findings of [63].
In the following plots, we compare the TROY results with those of the LL approximation, obtained with the low-energy parameters from the on-shell approach. Again, we use a small Gaussian source radius, $r_{0}=1$ fm, to enhance the effect of the strong interaction in the correlation function.
Figure 3: $\bar{D}^{0}p$ correlation function as function of the relative momentum. We use TROY with the solution of the off-shell $T$-matrix using the ZR approximation potential for the kernel. We use a Gaussian source radius of $r_{0}=1$ fm.
[…]
Figure 4: $D^{-}p$ correlation function as function of the relative momentum. We use TROY with the solution of the off-shell $T$-matrix using the ZR approximation potential for the kernel. We use a Gaussian source radius of $r_{0}=1$ fm.
[…]
The comparison with the LL approximation (short-dashed red line) shows some quantitative differences. While the TROY result remains always below the Coulomb only correlation function, and stays below one, the LL result remains always above one, and crosses the Coulomb only correlation function around $k=120$ MeV.
Figure 5: $D^{0}p$ correlation function as function of the relative momentum. We use TROY with the solution of the off-shell $T$-matrix using the ZR approximation potential for the kernel. We use a Gaussian source radius of $r_{0}=1$ fm.
[…]
Figure 6: $D^{+}p$ correlation function as function of the relative momentum. We use TROY with the solution of the off-shell $T$-matrix using the ZR approximation potential for the kernel. We use a Gaussian source radius of $r_{0}=1$ fm.
In Table 3 we provide a summary of the $s$-wave scattering lengths for all channels, in both on-shell and off-shell approximations of the ZR interaction.
Channel $a_{0}$ (fm) ZR TROY $a_{0}$ (fm) ZR on shell
[…]
Table 3: $s-$wave scattering length for $DN$ and $\bar{D}N$ channels in the ZR approximation using the off-shell unitarization implemented in TROY, and the on-shell factorization.
III.3 Coupled channels in the $D^{+}p$ and $D^{0}p$ correlation functions
In this section, we comment on the different channel contributions to the correlation function of $DN$ pairs. We first recall that the thermal weight $w_{i}$ associated to each channel coupling to the measured $DN$ pair has been obtained following the prescription of Eqs. (21) and (22). For further reference, the resulting values are provided in Table 4 for $D^{+}p$ pairs (left side) and $D^{0}p$ pairs (right side) and their related coupled channels.
Pair, $i$ $w_{i}$ Pair, $i$ $w_{i}$
$\pi^{+}\Lambda_{c}^{+}$ 2.20 $\pi^{0}\Lambda_{c}^{+}$ 2.16
$\pi^{0}\Sigma_{c}^{++}$ 0.92 $\pi^{0}\Sigma_{c}^{+}$ 0.90
$\pi^{+}\Sigma_{c}^{+}$ 0.91 $\pi^{+}\Sigma_{c}^{0}$ 0.90
$D^{+}p$ 1.00 $\pi^{-}\Sigma_{c}^{++}$ 0.90
$K^{+}\Xi_{c}^{+}$ 0.29 $D^{0}p$ 1.00
$\eta\Sigma_{c}^{++}$ 0.26 $D^{+}n$ 0.97
$K^{+}\Xi^{\prime+}_{c}$ 0.16 $\eta\Lambda_{c}^{+}$ 0.60
$D_{s}^{+}\Sigma^{+}$ 0.19 $K^{+}\Xi_{c}^{0}$ 0.28
$\eta^{\prime}\Sigma_{c}^{++}$ 0.04 $K^{0}\Xi_{c}^{+}$ 0.28
$\eta\Sigma_{c}^{+}$ 0.25
$K^{+}\Xi_{c}^{\prime 0}$ 0.16
$K^{0}\Xi_{c}^{\prime+}$ 0.16
$D_{s}^{+}\Lambda^{0}$ 0.26
$D_{s}^{+}\Sigma^{0}$ 0.18
$\eta^{\prime}\Lambda_{c}^{+}$ 0.10
$\eta^{\prime}\Sigma_{c}^{+}$ 0.04
Table 4: Thermal weights, as computed by the prescription described in Sec. II.3.2, of the $D^{+}p$ (left side) and $D^{0}p$ (right side) cases and their coupled channels.
The decomposition of the $D^{+}p$ and the $D^{0}p$ correlation functions for a source size of $r_{0}=1$ fm is shown in Fig. 7.
The left panel of this figure focuses on the $D^{+}p$ correlation function. The main contribution comes from the elastic channel, but the remaining ones (added one by one from the bottom to the top) contribute considerably. Notice that this channel is only affected by the $I=1$ sector of the $DN$ interaction, where the dynamically generated $\Sigma_{c}(2800)$ appears, a state that has also been experimentally seen in $\pi\Lambda_{c}^{+}$ spectra [62]. According to Ref. [63] this resonance couples very strongly to $DN$. It also has a non-negligible coupling to the $\pi\Lambda_{c}$ and $\pi\Sigma_{c}$ channels, the size of the former being twice that of the latter. These two channels also have the largest thermal weights, since their thresholds are located below the $DN$ one. The combined effect of the weight and the coupling to the $\Sigma_{c}(2800)$ explains the large contribution of the $\pi\Lambda_{c}$ channel (difference between the solid purple and dashed red lines), followed by that of the $\pi\Sigma_{c}$ channels (difference between the dotted blue and dashed red lines; and between the dashed green and dotted blue lines). The rest of the channels have smaller weights, so their effect is negligible. We conclude that a measurement of the $D^{+}p$ correlation function would be very sensitive to the presence of the $\Sigma_{c}(2800)$ and would bring very valuable information to constrain the models of the $DN$ interaction.
In the right panel, we plot the $D^{0}p$ case and its coupled channels, one by one, from the bottom to the top. Here we observe a large effect of the $\pi\Lambda_{c}$ channel, for the same reason as before, that is, because of its coupling to the $I=1$ $\Sigma_{c}(2800)$ resonance and the larger value of the thermal weight. For the $\pi\Sigma_{c}$ case we observe a moderate effect from the $\pi^{+}\Sigma_{c}^{0}$ (difference between the solid green and dashed blue lines) and $\pi^{-}\Sigma_{c}^{++}$ (difference between the solid orange and solid green lines) channels, which are coupled to the $D^{0}p$ state in both the $I=0$ and $I=1$ channels, dominated respectively, by the $\Lambda_{c}(2595)$ and the $\Sigma_{c}(2800)$ resonances. Interestingly, the effect of the $\pi^{0}\Sigma_{c}^{+}$ (difference between the dashed red and blue lines) is almost zero. We note that this state only couples to the $D^{0}p$ via the $I=0$ channel (the $I=1$ has zero Clebsch-Gordan coefficient). We therefore conclude that the relevance of a channel in the $D^{0}p$ correlation function is driven by its coupling to the $\Sigma_{c}(2800)$ resonance, which lies close to the kinematical region of interest. Finally, we observe that the largest coupled channel contribution is due to the $D^{+}n$ channel, which has a threshold very close to the $D^{0}p$ one, and therefore a similar thermal weight. This contribution is enhanced by the fact that the $DN$ coupling to the $\Sigma_{c}(2800)$ resonance is about an order of magnitude larger than that of the $\pi\Lambda_{c}$ and $\pi\Sigma_{c}$ states [63]. The channels with thresholds above the $D^{+}n$ one provide a negligible contribution to the $D^{0}p$ correlation function.
Figure 7: Correlation function of $DN$ for a Gaussian source with $r_{0}=1$ fm. Left panel: $D^{+}p$ with only strong interaction. Right panel: $D^{0}p$. The lines are to be read from bottom to the top when we add additional channels up to complete all of them (“all channels”)
III.4 Dependence on the source size and comparison with data
In Fig. 8 we plot the $D^{-}p$ correlation function using TROY for the meson-baryon interaction in the ZR approximation. It represents an overall (Coulomb plus strong) attractive interaction, which cannot fully explain the current experimental data by ALICE [27]. We have used a Gaussian radius between 0.8 fm and 1.0 fm, relevant for pp collisions, and present the results with a blue band. In Ref. [27] only the parametrizations by Haidenbauer—with an attractive interaction in the strong sector given by an additional strength of the scalar exchange—and Yamaguchi, with a bound state, show a reasonable agreement with the data. In the Yamaguchi case, as can be seen in our Fig. 2, we obtain a result consistent with the presence of the bound state. We have not been able to reproduce the curve shown in the ALICE paper [27] for this model, unless one naively assumes $a_{D^{-}p}\simeq a_{\bar{D}N,I=0}$.
Figure 8: $D^{-}p$ correlation function using TROY solution of the ZR approximation. The Gaussian radii is varied between $r_{0}=(0.9\pm 0.1)$ fm, to cover small systems as those measured by ALICE in high-multiplicity pp collisions.
Figure 9: $D^{0}p$ (top) and $\bar{D}^{0}p$ (bottom) correlation functions using TROY solution of the ZR approximation with thermal weights. The Gaussian radii are taken to be 1 fm, 3 fm, and 5 fm to cover larger systems as those measured by the STAR collaboration in these sectors
In Fig. 9 we present the results of the $D^{0}p$ and $\bar{D}^{0}p$ channels using the full coupled-channel TROY formalism with thermal weights. The Gaussian radii are chosen to be $r_{0}=1$ fm, 3 fm, and 5 fm. The latter is used to mimic the source sizes expected in Au+Au collisions measured by the STAR collaboration. The signal for these large radii is rather weak and suppressed, but our result is roughly consistent with the preliminary data shown in Ref. [64]. However, in that reference the $D^{0}p$ and $\bar{D}^{0}p$ cases are summed up to a single correlation function. Since the interaction is rather different for each of these sectors, we doubt this is a sensible comparison and urge our experimental colleagues to provide each channel separately.
IV Summary
In this work we have studied the femtoscopic correlation function of $D^{+}p,D^{0}p,D^{-}p$ and $\bar{D}^{0}p$ pairs applying the Koonin-Pratt equation. The meson-baryon interaction has been extracted from an effective field theory at low energy, working in the zero-range approximation. Non perturbative $T$-matrix equations have been implemented both in the on-shell factorization approach, as well as in the off-shell case using the TROY formalism. These two methods provide different approximations for the $l=0$ wave function of the meson-baryon pair, which eventually are used in the Koonin-Pratt formula. Coulomb effects have been also incorporated in the appropriate channels.
The on-shell calculation was used to extract low-energy parameters of the interaction which were employed in the Lednický-Lyuboshitz approximation of the correlation function. We remind the reader that this approximation is based on the asymptotic behavior of the wave function of the direct channel, thereby neglecting the effect of the inelastic components. For the $D^{+}p$ case we found that different results in the literature (with the exception of the SU(8) calculation) provide a similar correlation function reflecting a repulsive strong interaction. For the $D^{-}p$ pair we also find a consistency among the different results despite the different values of the scattering lengths, with a clear difference of the Haidenbauer’s calculation including an enhanced $\sigma$ exchange which encodes a rather strong attraction. This result was close to the experimental points as shown in Ref. [8]. We stress that in our version of the LL formula, the Yamaguchi result does not approach the data, but shows the typical effect of a bound state, in accordance with the theoretical expectations.
The full TROY calculation incorporates all the inelastic channels and obtains the complete spatial dependence of the wave function (not just the asymptotic part) out of the off-shell $T$-matrix elements. Moreover, we have estimated the thermal weights for all channels taking into account the masses of the produced pairs in a system at finite temperature for conditions given at high-multiplicity pp collisions at $\sqrt{s}=13$ TeV in the LHC. For the $D^{+}p$ case we found in our model that the direct channel is roughly compatible with the LL approximation, but the effect of the coupled channels is non negligible, increasing the correlation function between 30-50% at low momentum. We provide this correlation function for a small source radius as a prediction for future measurements of this system in the ALICE experiment, in which we foresee that the LL approach will not be enough. For the $D^{-}p$ case, the effect of the only coupled channel $(\bar{D}^{0}n)$ is found to minimal with respect to the direct one, but there are differences between the LL and TROY calculations of the correlation function, the latter maintaining values above one for all momenta, while the former approaches one from below. We find no consistency with the experimental data for $D^{+}p$ pairs measured by ALICE in pp collisions. Finally, we have computed the correlation function for neutral channels, $D^{0}p$ and $\bar{D}^{0}p$, motivated by the preliminary results by the STAR collaboration. Since these correspond to Au+Au collisions we scanned large Gaussian radii between $r_{0}=1$ fm and $r_{0}=5$ fm. We have not quantitatively compared with preliminary data since they are given as the sum of the two channels, and we cannot really assess the different nature of the interaction in the $C=1$ and $C=-1$ sectors.
To further test the results of this paper, it would be vital to compare the $D^{+}p$ correlation function in small systems like pp collisions, for which we present a prediction but no experimental results have been provided yet by the ALICE experiment. This channel could be an excellent way to study the effect of the $\Sigma_{c}(2800)$ resonance. In addition, it would be interesting to compare with STAR results in larger systems, like the one shown in [64], but after the data of the two charm sectors are given separately. While this would lead to an increase of the statistical uncertainties, it would make the comparison more sensible, since in our model the $D^{0}p$ and $\bar{D}^{0}p$ interactions are rather different, namely, strongly attractive and repulsive, respectively.
Acknowledgements.
We acknowledge support from the following national agencies: Unidades de Excelencia María de Maeztu CEX2020-001058-M and CEX2024-001451-M, and Projects No. PID2022-139427NB-I00 and No.PID2023-147112NB-C21, funded by MICIU/AEI/10.13039/501100011033 (Spain); Contract 2021 SGR 171 by the Generalitat de Catalunya; CRC-TR 211 ’Strong-interaction matter under extreme conditions’, Project No. 315477589 - TRR 211 by the Deutsche Forschungsgemeinshaft; Grant No. 402942/2024 by the Brazilian CNPq (National Council for Scientific and Technological Development) and Grant CIPROM 2023/59 of Generalitat Valenciana.